### v3

@@ Appendix A Details of DexManip tasks
  • Tool: This dataset consists of 200 rollouts that continues after the nail is driven into the board, so that the hammer stays stable in the air ($T=100$). This is the original setting proposed in Rajeswaran et al. (2018a).
@@
This task is considered i) successful if at last time step (note that during evaluation, the policy will always run for 100 time steps, similar to the Tool dataset), the Euclidean distance between the final nail position and the goal nail position is smaller than 0.01m (Han et al., 2023), and ii) safe if the final distance from the hammer to the goal nail position is smaller than 0.3m, indicating that the robot hand neither drifts away nor throws away the hammer.
@@
At each time step, if the Euclidean distance between the current cylinder position and the target position is smaller than 0.10m, then we have $\rho(t)=1$. This task is considered i) successful if $\sum_{t=1}^{T=100}\rho(t)>10$ (this criterion is the same as that used in Han et al. (2023)), and ii) safe if the robot hand is still near the target position (i.e., the distance is smaller than 0.5m) at final time step, indicating that the robot hand does not drifts away.
Appendix B Spectrum Clipping for Controlled System
@@
where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n},\mathrm{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ are the system state and control signal at time $t$, respectively. Similarly, $\{A,B\}\in\mathcal{S}_{A,B}\subset\{\mathbb{R}^{n\times n},\mathbb{R}^{n\times m}\}$, where $A,B$ are the system evolution matrix and the control matrix. We use $\mathcal{S}_{A,B}$ to denote the solution space of the matrices $A$ and $B$ that describes the system evolution of Equation. 5.
Now, suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{u}_{1},\mathrm{x}_{2},\mathrm{u}_{2},\cdots,\mathrm{x}_{T}]$. Let $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, and $U=[\mathrm{u}_{1},\mathrm{u}_{2},\cdots,\mathrm{u}_{T-1}]\in\mathbb{R}^{m\times T-1}$, then to learn the linear system from the dataset $D$ is to minimize the regression objective as follow:
$A,B=\arg\min_{\mathcal{S}_{A,B}}||Y-AX-BU||_{F}^{2},$ (6)
Similarly, optimizing Equation. 6 does not guarantee stability. Here, the stability constraint for the controlled system is similar to that of an autonomous system, with the exception that the solution space changes to $\mathbb{S}_{A,B}=\{\{A,B\}\in\mathcal{S}_{A,B}|\max\{\lambda_{i}(A)\}^{n}_{i=1}<=1\}$. Therefore, the objective in Equation. 6 turns into
$A,B=\arg\min_{\mathbb{S}_{A,B}}||Y-AX-BU||_{F}^{2},$ (7)
Learning Stable Linear Systems from Data with Control Inputs:
To find the approximately optimal solution to Eqn. 7, to We first compute $A_{ls}$ and $B_{ls}$ by optimizing the least square regression (Eqn. 6). Then we clip matrix $A_{ls}$ to $\bar{A_{sc}}$ as described in Section. 3, while keeping $B_{ls}$ unchanged.
Appendix C Qualitative performance on the Franka Panda dataset
@@
Appendix D Tradeoff between stability and expressivity
As mentioned in Sec. 3, we can also select the $\varepsilon$ values to ensure that the resulting linear system meets the criteria for well-known Lyapunov stability. Through the experiments conducted on UCSD, DTDB, DexManip, and Franka Panda datasets, we demonstrate that this strategy does not only guarantee the Lyapunov stability, but also achieve comparable performance, provided that an appropriate $\varepsilon$ is chosen, such as $\varepsilon=10^{-5}$. Note that we omit the computation time and memory usage results, as this strategy does not impact these metrics.
@@ Appendix D Tradeoff between stability and expressivity
Figure 12: Average reconstruction error evaluated at each time step percentile of the DTDB dataset.
Figure 13: Average Reconstruction error over each dataset. Solid lines represent median values, and shaded areas indicate quartile values.
and show the results in Fig. 13 and Fig. 13, respectively. We can observe that the Reconstruction error of SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ are not distinguishable. However, if the $\varepsilon$ is chosen larger, i.e., $\varepsilon=10^{-2}$, the error becomes significantly larger, primarily because the learned system is too conservative and lacks expressivity.
@@
For Franka Panda dataset, though SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ still perform similarly, we observe that $SC(\varepsilon=10^{-2})$ also performs similarly, and even performs the best when using the fewest data points for training (see Fig. 19). We conjecture that this is due to the control inputs and feedback controller. In addition, all these three performs comparably in the long-term prediction (Fig. 20).
@@
Turning black ($\downarrow$) Blowing up ($\downarrow$) Normal ($\uparrow$) Moving ($\uparrow$)
LS 6.67% 45.96% 47.37% 62.11%
SC($\varepsilon=0$) 14.39% 0% 85.61% 54.39%
SC($\varepsilon=10^{-5}$) 14.74% 0% 85.26% 54.04%
SC($\varepsilon=10^{-2}$) 100.$0$% 0%
Table 4: Quantitative ratios of the characterization.
@@
Table 6: Safety Rate for the learned policy with varying $\epsilon$ values.
@@
Figure 20: Distance to correct end effector position calculated at each time step.

### v4

@@ Appendix B Details of DexManip tasks
  • Tool: This dataset consists of 200 rollouts that continues after the nail is driven into the board, so that the hammer stays stable in the air ($T=100$). This is the original setting proposed in [47].
@@
This task is considered i) successful if at last time step (note that during evaluation, the policy will always run for 100 time steps, similar to the Tool dataset), the Euclidean distance between the final nail position and the goal nail position is smaller than 0.01m [13], and ii) safe if the final distance from the hammer to the goal nail position is smaller than 0.3m, indicating that the robot hand neither drifts away nor throws away the hammer.
@@
At each time step, if the Euclidean distance between the current cylinder position and the target position is smaller than 0.10m, then we have $\rho(t)=1$. This task is considered i) successful if $\sum_{t=1}^{T=100}\rho(t)>10$ (this criterion is the same as that used in [13]), and ii) safe if the robot hand is still near the target position (i.e., the distance is smaller than 0.5m) at final time step, indicating that the robot hand does not drifts away.
Appendix C Spectrum Clipping for Controlled System
@@
where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n},\mathrm{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ are the system state and control signal at time $t$, respectively. Similarly, $A\in\mathcal{S}_{A}$ is the unknown system matrix, and $\mathcal{S}_{A}\subset\mathbb{R}^{n\times n}$ is the subspace that contains stable matrices: $\mathcal{S}_{A}\triangleq\{A\in\mathbb{R}^{n\times n}|\max\{|\lambda_{i}(A)|\}^{n}_{i=1}<=1\}$, and $B\in\mathcal{S}_{B}$ is the unknown control matrix, and $\mathcal{S}_{B}\subset\mathbb{R}^{n\times m}$.
Unconstrained Learning: Now, suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{u}_{1},\mathrm{x}_{2},\mathrm{u}_{2},\cdots,\mathrm{x}_{T}]$. Let $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, and $U=[\mathrm{u}_{1},\mathrm{u}_{2},\cdots,\mathrm{u}_{T-1}]\in\mathbb{R}^{m\times T-1}$, then to learn the linear system from the dataset $D$ is to minimize the similar regression objective:
$\hat{A}_{LS},\hat{B}=\arg\min_{\{\Theta,\Gamma\}}||Y-AX-BU||_{F}^{2},$ (9)
where $\Theta\in\mathbb{R}^{n\times n}$ and $\Gamma\in\mathcal{S}_{B}$ are both the learnable parameters, and $||\cdot||_{F}$ is the matrix Frobenius norm. Similarly, optimizing Eq. (9) does not guarantee stability.
Learning Stable LDS with Control Inputs: To this end, we can rewrite the optimization problem Eq. (9) to include stability constraints:
$\hat{A}_{S},\hat{B}=\arg\min_{\{\Theta\in\mathcal{S}_{A},\Gamma\}}||Y-AX-BU||_{F}^{2},$ (10)
To find the approximate solution to Eq. (10), we first compute $\hat{A}_{LS}$ and $\hat{B}$ by optimizing the least square regression (Eq. (9)). Then we apply SC method to $\hat{A}_{LS}$ and obtain $\hat{A}_{SC}$ as described in Section. 3, while keeping $\hat{B}$ unchanged.
Appendix D Qualitative performance on the Franka Panda dataset
@@
Appendix E Tradeoff between stability and expressivity
As mentioned in Section. 3, we can also select the $\varepsilon$ values to ensure that the resulting linear system meets the criteria for well-known Lyapunov stability. Through the experiments conducted on UCSD, DTDB, DexManip, and Franka Panda datasets, we demonstrate that this strategy does not only guarantee the Lyapunov stability, but also achieve comparable performance, provided that an appropriate $\varepsilon$ is chosen, such as $\varepsilon=10^{-5}$. Note that we omit the computation time and memory usage results, as this strategy does not impact these metrics.
@@ Appendix E Tradeoff between stability and expressivity
Figure 10: Left figure: average reconstruction error evaluated on different subspace dimensions of the UCSD dataset. Solid lines represent median values, and shaded areas indicate quartile values. Right figure: Average reconstruction error evaluated at each time step percentile of the DTDB dataset. Solid lines represent median values, and shaded areas indicate quartile values.
@@
and show the results in Fig. 10. We can observe that the Reconstruction error of SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ are not distinguishable. However, if the $\varepsilon$ is chosen larger, i.e., $\varepsilon=10^{-2}$, the error becomes significantly larger, primarily because the learned system is too conservative and lacks expressivity.
@@
For Franka Panda dataset, though SC($\varepsilon=0$) and $SC(\varepsilon=10^{-5})$ still perform similarly, we observe that $SC(\varepsilon=10^{-2})$ also performs similarly, and even performs the best when using the fewest data points for training (see Fig. 17). We conjecture that this is due to the control inputs and feedback controller. In addition, all these three performs comparably in the long-term prediction (Table. 10).
@@
Average NIQE ($F=500$) ($\downarrow$) Average NIQE ($F=3000$) ($\downarrow$) Moving ($\uparrow$)
LS 904.6265 6355.5817 62.11%
SC($\varepsilon=0$) 18.8769 1075.0140 54.39%
SC($\varepsilon=10^{-5}$) 18.8769 1075.0139 54.04%
SC($\varepsilon=10^{-2}$) 706.0687 inf 0%
[…]
Table 7: Quantitative metrics evaluating the predictions with varying $\varepsilon$ values on the DTDB dataset.
@@
Table 9: Safety Rate for the learned policy with varying $\varepsilon$ values.
@@
LS SC($\varepsilon=0$) SC($\varepsilon=10^{-5}$) SC($\varepsilon=10^{-2}$)
$d=100$ $0.0748\pm 0.1218$ $0.0693\pm 0.1370$ $0.0686\pm 0.1344$ $0.0193\pm 0.0074$
$d=2000$ $0.0207\pm 0.0173$ $0.0183\pm 0.0138$ $0.0182\pm 0.0132$ $0.0175\pm 0.0062$
$d=3000$ $0.0164\pm 0.0083$ $0.0156\pm 0.0057$ $0.0156\pm 0.0057$ $0.0176\pm 0.0062$
Table 10: Average prediction error of the end-effector’s position (unit: $m$).